OpenStudy (anonymous):

In a right triangle, find the length of the side not given. b=2, c=sqrt13, a=__

OpenStudy (anonymous):

which one is the hypotenuse?

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

c is usually hypotenuse

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

ya thats a really bad assumption to make always

OpenStudy (anonymous):

if its not given then thats the only way we can solve it. unless there was a picture associated with the question its safe to assume that

OpenStudy (anonymous):

\[c ^{2}=a ^{2}+b ^{2}\]hence \[13=a ^{2}+4\] hence a=\[\sqrt{9}\]=3

hero (hero):

\[\color{green} {\huge \checkmark} \]